OpenStudy (anonymous):

Compute 4659¯0−214¯00. Round the answer appropriately. Express your answer numerically using the proper number of significant figures.

OpenStudy (anonymous):

my answer is 4450.0. however, it's incorrect.

OpenStudy (anonymous):

can someone tell me what i did wrong?

OpenStudy (anonymous):

I'm not sure what the question is. It looks like it might be \(4659.0-214.00\) if that is the case, it is just subtraction. I'd say to check your arithmetic with a calculator.

OpenStudy (anonymous):

well, i did do that but still got it wrong

OpenStudy (anonymous):

like before

OpenStudy (anonymous):

My point is you subtracted incorrectly. You need to line up the decimals to add or subtract. 9-5=5; 5-1=4; 6-2=4; 4-0=4; \(~4659.0\) \(\underline{-214.00}\) \(~4445.00\) As for the format of the answer, I would guess that it should be 1 decimal place since that is the smallest number of decimal places in the question.